# Flexible inference in heterogeneous and attributed multilayer networks

### v1

@@ Abstract
Networked datasets are often enriched by different types of information about individual nodes or edges. However, most existing methods for analyzing such datasets struggle to handle the complexity of heterogeneous data, often requiring substantial model-specific analysis. In this paper, we develop a probabilistic generative model to perform inference in multilayer networks with arbitrary types of information. Our approach employs a Bayesian framework combined with the Laplace matching technique to ease interpretation of inferred parameters. Furthermore, the algorithmic implementation relies on automatic differentiation, avoiding the need for explicit derivations. This makes our model scalable and flexible to adapt to any combination of input data. We demonstrate the effectiveness of our method in detecting overlapping community structures and performing various prediction tasks on heterogeneous multilayer data, where nodes and edges have different types of attributes. Additionally, we showcase its ability to unveil a variety of patterns in a social support network among villagers in rural India by effectively utilizing all input information in a meaningful way.
@@ Introduction
The analysis of attributed multilayer networks has primarily been tackled using techniques like matrix factorization [2, 3], network embedding [4, 5], and deep learning [6, 7, 8, 9]. However, in this work, we focus on a less investigated methodology that involves probabilistic generative models [10]. Unlike the aforementioned approaches, these methods provide a principled and flexible framework to incorporate prior knowledge and specific assumptions, while also accounting for the inherent uncertainty present in real-world data [11]. Moreover, they can be applied to perform inference on networks, including tasks such as predictions or the detection of statistically meaningful network structures. The latter task is commonly referred to as community detection problem, and is relevant in many applications [12].
Our goal is to develop a probabilistic generative model that can flexibly adapt to any attributed multilayer network, regardless of the type of information encoded in the data. Acting as a “black box”, our method can enable practitioners to automatically analyze various datasets, without the need to deal with mathematical details or new derivations. This approach aligns with some practices in the machine learning community, where principled black box methodologies have been introduced to simplify the inference of latent variables in arbitrary models [13, 14]. In this context, more specific probabilistic methods have been developed to address the challenge of performing inference on heterogeneous data [15, 16]. However, these techniques are tailored for tabular data and do not provide a general solution to adapt them to network data.
@@ Methods
Fig. 1: Input data and graphical model representation. (A) The attributed multilayer network is represented by the interactions $A_{ij}^{\ell}$ and the node attributes $X_{ix}$. (B) PIHAM describes the observed data through a set of latent variables $\boldsymbol{\Theta}=(\boldsymbol{U},\boldsymbol{V},\boldsymbol{W},\boldsymbol{H})$. $\boldsymbol{U_{i}}$ and $\boldsymbol{V_{i}}$ respectively depict the out-going and in-coming communities of node $i$; $\boldsymbol{W^{\ell}}$ is the affinity matrix associated to the layer $\ell$ and characterizes the edge density between different community pairs in the given layer; $\boldsymbol{H_{\cdot x}}$ is a $K$-dimensional vector that explains how an attribute $x$ is distributed among the $K$ communities. All latent variables are independent and normally distributed, and $f(\cdot)$ and $g(\cdot)$ are transformation functions to ensure that the expected values $\lambda_{ij}^{\ell}$ and $\pi_{ix}$ belong to the correct parameter space for the various distribution types.
@@
where the latent variables $U_{ik}$ and $V_{jq}$ denote the entries of $K$-dimensional vectors $\boldsymbol{U_{i}}$ and $\boldsymbol{V_{i}}$, which respectively represent the out-going and in-coming communities of node $i$. In undirected networks, we set $\boldsymbol{U}=\boldsymbol{V}$. Moreover, each layer $\ell$ is associated with an affinity matrix $\boldsymbol{W^{\ell}}$ of dimension $K\times K$, which characterizes the edge density between different community pairs in the given layer $\ell$. This setup allows having diverse structural patterns in each layer, including arbitrarily mixtures of assortative, disassortative and core-periphery structures.
@@
where $\boldsymbol{H}$ is a $K\times P$-dimensional community-covariate matrix, explaining how an attribute $x$ is distributed among the $K$ communities. For instance, if we consider income as node metadata and expect communities to group nodes with similar income values, then the column vector $\boldsymbol{H_{\cdot x}}$ describes how income varies across groups. It is important to observe that when the attribute $x$ is categorical, the expression in Eq. 5 becomes more complex because it must consider the total number of attribute categories $Z$. We provide additional details in the Supporting Information.
@@
In Fig. 1, we illustrate the input data and the graphical model representation of our approach.
@@
PIHAM aims at estimating the posterior distributions of the latent variables, as outlined in General framework. More precisely, this equation can be reformulated as:
@@
LA involves a second-order Taylor expansion around the Maximum A Posteriori estimate (MAP) of the right-hand side of Posterior inference. We compute this estimate using Automatic Differentiation (AD), a gradient-based method that, in our implementation, employs the Adam optimizer to iteratively evaluate derivatives of the log-posterior. The MAP estimate found with AD also constitutes the mean $\hat{\boldsymbol{\mu}}^{\boldsymbol{\theta}}$ of $P(\boldsymbol{\theta}\,|\,\boldsymbol{A},\boldsymbol{X})$. To go beyond point estimates and quantify uncertainty, one can further estimate the covariance matrix $\hat{\boldsymbol{\Sigma}}^{\boldsymbol{\theta}}$, which is given by the inverted Hessian around the MAP:
@@ Results
We demonstrate our method on both synthetic and real-world datasets, presenting a comprehensive analysis through quantitative and qualitative findings. Further explanations about the data generation and pre-processing procedures can be found in the Supporting Information, which also includes additional results. The settings used to run our experiments and the choice of the hyperparameters are also described in the Supporting Information. The code implementation of PIHAM is accessible at: https://github.com/mcontisc/PIHAM.
@@
We analyze attributed multilayer networks with $L=3$ heterogeneous layers: one with binary interactions, one with nonnegative discrete weights, and one with real values. In addition, each node is associated with three covariates: one categorical with $Z=4$ categories, one representing nonnegative discrete values, and one involving real values. To generate these networks, we initially draw the latent variables $\boldsymbol{\Theta}=(\boldsymbol{U},\boldsymbol{V},\boldsymbol{W},\boldsymbol{H})$ from Gaussian distributions with specified hyperparameters. Subsequently, we generate $\boldsymbol{A}$ and $\boldsymbol{X}$ according to the data types, following Modelling the network structure and Modelling the node metadata. Our analysis spans networks with varying number of nodes $N\in\{100,200,\dots,1000\}$ and diverse number of overlapping communities $K\in\{3,4,5\}$. Additional details on the generation process can be found in the Supporting Information.
@@
Computing the Overlap for many communities can be computationally expensive due to the need to calculate all pairwise combinations. As an alternative solution, we suggest utilizing the $L_{2}$-barycenter distribution, which essentially represents a weighted average of the node-community distributions [32, 33]. We show the barycenter distributions in gray in the second row of Fig. 3. This approach allows focusing on a single distribution per node, instead of $K$ different ones. To quantify this distribution, we calculate its variance ($\sigma^{2}$), where higher values indicate nodes with harder memberships, as the barycenter is more spread due to the individual distributions being more distant from each other. Conversely, lower variance suggests more overlap among the distributions, indicating a more mixed-membership scenario. We observe that $\sigma^{2}$ increases as we decrease the degree of mixed-membership, a trend consistent with that of the Overlap. Further details on the barycenter distribution and the metrics are provided in the Supporting Information.
@@
We now turn our attention to the analysis of a real-world dataset describing a social support network within a village in Tamil Nadu, India, referred to as “Alakāpuram” [34, 35]. The data were collected in 2013 through surveys, in which adult residents were asked to nominate individuals who provided various types of support, such as running errands, offering advice, and lending cash or household items. Additionally, several attributes were gathered, encompassing information like gender, age, and caste, among others. The pre-processing of the dataset is described in the Supporting Information. The resulting heterogeneous attributed multilayer network comprises $N=419$ nodes, $L=7$ layers, and $P=3$ node attributes. The initial six layers depict directed binary social support interactions among individuals, with average degree ranging from $1.8$ to $4.2$. The seventh, instead, contains information that is proportional to the geographical distance between individuals’ households. The adjacency tensor is then represented as $\boldsymbol{A}=\{\boldsymbol{A^{\ell}}\in\{0,1\}^{N\times N}\,\forall\ell\in[1,6],\boldsymbol{A^{7}}\in\mathbb{R_{+}}^{N\times N}\}$. As node covariates, we consider the caste attribute with $Z_{caste}=14$ categories, the religion attribute with $Z_{religion}=3$ categories, and the attribute representing the years of education, that is $\boldsymbol{X_{\cdot 3}}\in\mathbb{N}_{0}^{N}$. Ethnographic work and earlier analyses [35, 36] suggest that these attributes play an important role in how villagers relate to one another, with certain relationships being more strongly structured by these identities than others.
Inference and prediction performance.
@@
We describe the likelihood of the real-world heterogeneous attributed multilayer network according to Modelling the network structure and Modelling the node metadata, customized to suit the data types under examination. In particular, we employ Bernoulli distributions for the binary layers $[\boldsymbol{A^{\ell}}]_{\ell\in[1,6]}$ and Gaussian distributions for the distance layer $\boldsymbol{A^{7}}$. Moreover, we characterize the attributes caste $\boldsymbol{X_{\cdot 1}}$ and religion $\boldsymbol{X_{\cdot 2}}$ using Categorical distributions, and model the covariate $\boldsymbol{X_{\cdot 3}}$ with a Poisson distribution. The choice of the model hyperparameters and the algorithmic settings used in our experiments are described in the Supplementary Information.
@@
Similarly to many real-world datasets, we lack the information about the true parameters underlying the network, including the node memberships. Hence, to determine the number of communities $K$, we employ a $5$-fold cross-validation procedure for $K\in[1,10]$ and select the value that exhibits the optimal performance. Detailed results are displayed in Table S3 of the Supporting Information. We set $K=6$ as it achieves the best performance across the majority of prediction metrics. In fact, selecting a single metric to summarize and evaluate results in a heterogeneous setting is nontrivial, as discussed in the section Validation on heterogeneous data..
The results in Table S3 additionally validate PIHAM’s performance in inference tasks like edge and covariate prediction. Overall, our method demonstrates robust outcomes with the chosen fixed value of $K$ and consistently outperforms the baselines, which are omitted for brevity.
@@ Discussion
In this work, we have introduced PIHAM, a probabilistic generative model designed to perform inference in heterogeneous and attributed multilayer networks. A significant feature of our approach is its flexibility to accommodate any combination of the input data, made possible through the use of Laplace approximations and automatic differentiation methods, which avoid the need for explicit derivations. Furthermore, PIHAM employs a Bayesian framework, enabling the estimation of posterior distributions, rather than only providing point estimates for the parameters.
@@
When compared to other methods tailored for scenarios with only one type of attribute and interaction, PIHAM demonstrates comparable performance in prediction and community detection tasks, despite its broader formulation. Moreover, our approach significantly outperforms baseline metrics in more complex settings characterized by various attribute and interaction types, where existing methods for comparison are lacking. Furthermore, PIHAM employs the Laplace matching technique, offering a theoretically grounded approach to map posterior distributions to various desired domains, facilitating interpretation.
@@
While PIHAM constitutes a principled and flexible method to analyze heterogeneous and attributed multilayer networks, several questions remain unanswered. For example, determining the most appropriate metric for summarizing prediction performance in heterogeneous scenarios, where information spans different spaces, is not straightforward. This aspect also influences the selection of the optimal model during cross-validation procedures. While we have provided explanations for our choices, we acknowledge that this remains an open question. Similarly, when dealing with many communities, summarizing posterior distributions becomes challenging due to computational constraints. We addressed this issue by employing $L_{2}$-barycenter distributions and proposing their variance to guide interpretation. Nevertheless, we believe there is still considerable room for improvement and exploration in this area. Our method could be further extended to accommodate distinct community-covariate contributions by integrating two separate $\boldsymbol{H}$ matrices for both in-coming and out-going communities, respectively. This modification will offer clearer insights into how covariates influence the partitions, especially when discrepancies arise between in-coming and out-going communities. Additionally, it would be interesting to expand this framework to incorporate higher-order interactions, an emerging area that has shown relevance in describing real-world data [38].
@@
Fig. S5: Prediction and community detection performance on synthetic data. We analyze synthetic attributed multilayer networks with $N=500$ nodes, $L=2$ layers (one being assortative and the other disassortative), a categorical attribute with $Z=3$ categories, $K=3$ overlapping communities, and increasing average degrees $\langle{k}\rangle\in\{10,15,20,\dots,50\}$. The results represent averages and confidence intervals over $20$ independent samples. For prediction tasks, we employ a $5$-fold cross-validation procedure. The evaluation metrics include (A) the AUC for edge prediction, with a baseline of $0.5$ corresponding to random choice, and (B) accuracy for covariate prediction. Here, MRF represents a baseline given by the predictions obtained from the maximum frequency in the training set, while Random denotes the uniform random probability over $Z$. (C) Community detection performance is assessed using CS. As inferred point estimates, we consider both the mean of transformed Dirichlet posterior distributions and the $\operatorname{softmax}$ transformation of $\hat{\boldsymbol{\mu}}^{\boldsymbol{\theta}}$. Overall, PIHAM exhibits comparable performance to MTCOV across all tasks despite its broader framework, especially in scenarios involving denser networks.

### v2

@@ Abstract
Networked datasets can be enriched by different types of information about individual nodes or edges. However, most existing methods for analyzing such datasets struggle to handle the complexity of heterogeneous data, often requiring substantial model-specific analysis. In this paper, we develop a probabilistic generative model to perform inference in multilayer networks with arbitrary types of information. Our approach employs a Bayesian framework combined with the Laplace matching technique to ease interpretation of inferred parameters. Furthermore, the algorithmic implementation relies on automatic differentiation, avoiding the need for explicit derivations. This makes our model scalable and flexible to adapt to any combination of input data. We demonstrate the effectiveness of our method in detecting overlapping community structures and performing various prediction tasks on heterogeneous multilayer data, where nodes and edges have different types of attributes. Additionally, we showcase its ability to unveil a variety of patterns in a social support network among villagers in rural India by effectively utilizing all input information in a meaningful way.
@@ Introduction
It is important to treat these systems as multilayer networks, rather than reducing them to a single layer (e.g., through aggregation), to avoid losing valuable information. This is particularly significant when the properties of the entire system cannot be derived from a simple linear combination of the properties of each layer in isolation, or through other forms of aggregation into a single layer representation [2]. Relevant examples of this non-trivial behavior are observed in interbank markets, patent citation networks, time-dependent networks and brain networks  [3, 4, 5, 6]. A similar rationale applies to datasets with multiple types of edges and node attributes, as explored in this work. Approaches to multilayer networks that consider only one type of attribute – or none at all – risk losing critical information.
The analysis of attributed multilayer networks has primarily been tackled using techniques like matrix factorization [7, 8], network embedding [9, 10], and deep learning [11, 12, 13, 14]. While effective for learning low-dimensional node representations for inference tasks, these methods have significant limitations when applied to network data. First, the resulting representations lack inherent interpretability. Since their main focus is prediction tasks, these models do not impose constraints on the parameters to improve interpetability, such as non-negativity or shared parameters across layers. Instead, they often require arbritarily post-processing techniques (e.g., k-means clustering) for interpretation. Second, these methods prioritize node-level tasks (e.g., classification or clustering) and treat edge-level tasks, like link prediction, as secondary, relying on ad hoc mappings from nodes to edges. Lastly, they are typically parameter-heavy and require a substantial amount of training data, making them inefficient and overparametrized for real-world networks, which are typically sparse, label-scarce for supervised settings, and often provide only a single observed sample.
To overcome these limitations, we adopt a different approach based on probabilistic generative models [15]. Unlike the aforementioned methods, these models provide a principled and flexible framework that incorporates prior knowledge and specific assumptions, resulting in more interpretable representations. Importantly, they explicitly capture the dependencies between nodes and edges, rather than relying on ad hoc mappings. They also account for the inherent uncertainty present in real-world network data [16], providing a more robust and comprehensive understanding of this type of data. Furthermore, these models can be applied to perform various network tasks, such as edge and attribute prediction, detecting statistically meaningful network structures, and generating synthetic data.
Our goal is to develop a probabilistic generative model that can flexibly adapt to any attributed multilayer network, regardless of the type of information encoded in the data. Acting as a “black box”, our method can enable practitioners to automatically analyze various datasets, without the need to deal with mathematical details or new derivations. This approach aligns with some practices in the machine learning community, where black box methodologies have been introduced to simplify the inference of latent variables in arbitrary models [17, 18]. In this context, more specific probabilistic methods have been developed to address the challenge of performing inference on heterogeneous data [19, 20]. However, these techniques are tailored for tabular data and do not provide a general solution to adapt them to network data.
@@ Methods
Fig. 1: Input data and graphical model representation. (A) The attributed multilayer network is represented by the interactions $A_{ij}^{\ell}$ and the node attributes $X_{ix}$. (B) PIHAM describes the observed data through a set of latent variables $\boldsymbol{\Theta}=(\boldsymbol{U},\boldsymbol{V},\boldsymbol{W},\boldsymbol{H})$. $\boldsymbol{U_{i}}$ and $\boldsymbol{V_{i}}$ respectively depict the communities of node $i$ determined by the out-going and in-coming edges; $\boldsymbol{W^{\ell}}$ is the affinity matrix associated to the layer $\ell$ and characterizes the edge density between different community pairs in the given layer; $\boldsymbol{H_{\cdot x}}$ is a $K$-dimensional vector that explains how an attribute $x$ is distributed among the $K$ communities. All latent variables are independent and normally distributed, and $f(\cdot)$ and $g(\cdot)$ are transformation functions to ensure that the expected values $\lambda_{ij}^{\ell}$ and $\pi_{ix}$ belong to the correct parameter space for the various distribution types.
@@
where the latent variables $U_{ik}$ and $V_{jq}$ denote the entries of $K$-dimensional vectors $\boldsymbol{U_{i}}$ and $\boldsymbol{V_{i}}$, which respectively represent the communities of node $i$ determined by the out-going and in-coming edges. In undirected networks, we set $\boldsymbol{U}=\boldsymbol{V}$. Moreover, each layer $\ell$ is associated with an affinity matrix $\boldsymbol{W^{\ell}}$ of dimension $K\times K$, which characterizes the edge density between different community pairs in the given layer $\ell$. This setup allows having diverse structural patterns in each layer, including arbitrarily mixtures of assortative, disassortative and core-periphery structures.
@@
where $\boldsymbol{H}$ is a $K\times P$-dimensional community-covariate matrix, explaining how an attribute $x$ is distributed among the $K$ communities. For instance, if we consider income as node metadata and expect communities to group nodes with similar income values, then the column vector $\boldsymbol{H_{\cdot x}}$ describes how income varies across groups. It is important to observe that when the attribute $x$ is categorical, the expression in Eq. 5 becomes more complex because it must consider the total number of attribute categories $Z$. We provide additional details in Section S1 of the Supporting Information.
@@
PIHAM aims at estimating the posterior distributions of the latent variables, as outlined in General framework, where the normalization term is omitted as it does not depend on the parameters. More precisely, this equation can be reformulated as:
@@
LA involves a second-order Taylor expansion around the Maximum A Posteriori estimate (MAP) of the right-hand side of Posterior inference. We compute this estimate using Automatic Differentiation (AD), a gradient-based method that, in our implementation, employs the Adam optimizer to iteratively evaluate derivatives of the log-posterior. One might question the validity of this approximation in a potentially multimodal landscape. However, it is important to recognize that the existence of local minima is a problem independent of the approximation method employed. Nonetheless, in our model class – multilevel networks of log-convex likelihood functions connected by log-convex link functions – this issue is less likely to occur.
The MAP estimate found with AD also constitutes the mean $\hat{\boldsymbol{\mu}}^{\boldsymbol{\theta}}$ of $P(\boldsymbol{\theta}\,|\,\boldsymbol{A},\boldsymbol{X})$. To go beyond point estimates and quantify uncertainty, one can further estimate the covariance matrix $\hat{\boldsymbol{\Sigma}}^{\boldsymbol{\theta}}$, which is given by the inverted Hessian around the MAP:
@@ Results
We demonstrate our method on both synthetic and real-world datasets, presenting a comprehensive analysis through quantitative and qualitative findings. Further explanations about the data generation and pre-processing procedures can be found in the Supporting Information, which also includes additional results. The settings used to run our experiments and the choice of the hyperparameters are also described in the Supporting Information (see Section S2). The code implementation of PIHAM is accessible at: https://github.com/mcontisc/PIHAM.
@@
We analyze attributed multilayer networks with $L=3$ heterogeneous layers: one with binary interactions, one with nonnegative discrete weights, and one with real values. In addition, each node is associated with three covariates: one categorical with $Z=4$ categories, one representing nonnegative discrete values, and one involving real values. To generate these networks, we initially draw the latent variables $\boldsymbol{\Theta}=(\boldsymbol{U},\boldsymbol{V},\boldsymbol{W},\boldsymbol{H})$ from Gaussian distributions with specified hyperparameters. Subsequently, we generate $\boldsymbol{A}$ and $\boldsymbol{X}$ according to the data types, following Modelling the network structure and Modelling the node metadata. Our analysis spans networks with varying number of nodes $N\in\{100,200,\dots,1000\}$ and diverse number of overlapping communities $K\in\{3,4,5\}$. Additional details on the generation process can be found in Section S4 of the Supporting Information.
@@
Computing the Overlap for many communities can be computationally expensive due to the need to calculate all pairwise combinations. As an alternative solution, we suggest utilizing the $L_{2}$-barycenter distribution, which essentially represents a weighted average of the node-community distributions [36, 37]. We show the barycenter distributions in gray in the second row of Fig. 3. This approach allows focusing on a single distribution per node, instead of $K$ different ones. To quantify this distribution, we calculate its variance ($\sigma^{2}$), where higher values indicate nodes with harder memberships, as the barycenter is more spread due to the individual distributions being more distant from each other. Conversely, lower variance suggests more overlap among the distributions, indicating a more mixed-membership scenario. We observe that $\sigma^{2}$ increases as we decrease the degree of mixed-membership, a trend consistent with that of the Overlap. Further details on the barycenter distribution and the metrics are provided in Section S5 of the Supporting Information.
@@
We now turn our attention to the analysis of a real-world dataset describing a social support network within a village in Tamil Nadu, India, referred to as “Alakāpuram” [38, 39]. The data were collected in 2013 through surveys, in which adult residents were asked to nominate individuals who provided various types of support, such as running errands, offering advice, and lending cash or household items. Additionally, several attributes were gathered, encompassing information like gender, age, and caste, among others. The pre-processing of the dataset is described in Section S6 of the Supporting Information. The resulting heterogeneous attributed multilayer network comprises $N=419$ nodes, $L=7$ layers, and $P=3$ node attributes. The initial six layers depict directed binary social support interactions among individuals, with average degree ranging from $1.8$ to $4.2$. The seventh, instead, contains information that is proportional to the geographical distance between individuals’ households. The adjacency tensor is then represented as $\boldsymbol{A}=\{\boldsymbol{A^{\ell}}\in\{0,1\}^{N\times N}\,\forall\ell\in[1,6],\boldsymbol{A^{7}}\in\mathbb{R_{+}}^{N\times N}\}$. As node covariates, we consider the caste attribute with $Z_{caste}=14$ categories, the religion attribute with $Z_{religion}=3$ categories, and the attribute representing the years of education, that is $\boldsymbol{X_{\cdot 3}}\in\mathbb{N}_{0}^{N}$. Ethnographic work and earlier analyses [39, 40] suggest that these attributes play an important role in how villagers relate to one another, with certain relationships being more strongly structured by these identities than others.
Inference, prediction performance, and goodness of fit.
@@
We describe the likelihood of the real-world heterogeneous attributed multilayer network according to Modelling the network structure and Modelling the node metadata, customized to suit the data types under examination. In particular, we employ Bernoulli distributions for the binary layers $[\boldsymbol{A^{\ell}}]_{\ell\in[1,6]}$ and Gaussian distributions for the distance layer $\boldsymbol{A^{7}}$. Moreover, we characterize the attributes caste $\boldsymbol{X_{\cdot 1}}$ and religion $\boldsymbol{X_{\cdot 2}}$ using Categorical distributions, and model the covariate $\boldsymbol{X_{\cdot 3}}$ with a Poisson distribution. The choice of the model hyperparameters and the algorithmic settings used in our experiments are described in Section S2 of the Supplementary Information.
@@
Similarly to many real-world datasets, we lack the information about the true parameters underlying the network, including the node memberships. Hence, to determine the number of communities $K$, we employ a $5$-fold cross-validation procedure for $K\in[1,10]$ and select the value that exhibits the optimal performance. Detailed results are displayed in Table S3 of the Supporting Information. We set $K=6$ as it achieves the best performance across the majority of prediction metrics. In fact, selecting a single metric to summarize and evaluate results in a heterogeneous setting is nontrivial, as discussed in the section Validation on heterogeneous data.. The results in Table S3 additionally validate PIHAM’s performance in inference tasks like edge and covariate prediction. Overall, our method demonstrates robust outcomes with the chosen fixed value of $K$ and consistently outperforms the baselines, which are omitted for brevity.
We further evaluate our model’s goodness of fit through a posterior-predictive assessment [41, 42], comparing the input data to synthetic data generated by the fitted model. A well-fitted model should produce synthetic data that closely resemble the original input. To accurately assess performance, we test whether two samples from the posterior-predictive distribution are generally more, equally, or less distant from each other than a sample from the posterior-predictive distribution compared to the input data [41]. We measure the distance using different metrics depending on the data type, and the results are shown in Fig. S7. The discrepancies between synthetic data samples consistently exceed those between the observed data and synthetic samples, indicating that PIHAM  provides a good fit for the data.
@@ Discussion
In this work, we have introduced PIHAM, a probabilistic generative model designed to perform inference in heterogeneous and attributed multilayer networks. A significant feature of our approach is its flexibility to accommodate any combination of the input data, made possible through the use of Laplace approximations and automatic differentiation methods, which avoid the need for explicit derivations. However, it is important to note that having a method capable of handling complex network datasets does not automatically ensure the quality of the input data. For example, if only certain attributes are relevant for explaining the networked dataset, or if only specific layers contain valuable information for the task, adding unnecessary information could be detrimental. Practitioners must carefully assess which information is useful based on their specific objectives. Alternatively, model selection tests, such as the cross-validation routine demonstrated in the manuscript, can be used to determine the optimal combination of layers and attributes.
@@
When compared to other methods tailored for scenarios with only one type of attribute and interaction, PIHAM demonstrates comparable performance in prediction and community detection tasks, despite its broader formulation. Moreover, our approach significantly outperforms baseline metrics in more complex settings characterized by various attribute and interaction types, where existing methods for comparison are lacking. Additionally, PIHAM employs a Bayesian framework, enabling the estimation of posterior distributions, rather than only providing point estimates for the parameters. And, through the use of the Laplace matching technique, it maps these posterior distributions to various desired domains in a theoretically sound manner, facilitating interpretation.
@@
While PIHAM constitutes a principled and flexible method to analyze heterogeneous and attributed multilayer networks, several questions remain unanswered. For example, determining the most appropriate metric for summarizing prediction performance in heterogeneous scenarios, where information spans different spaces, is not straightforward. This aspect also influences the selection of the optimal model during cross-validation procedures. While we have provided explanations for our choices, we acknowledge that this remains an open question. Similarly, when dealing with many communities, summarizing posterior distributions becomes challenging due to computational constraints. We addressed this issue by employing $L_{2}$-barycenter distributions and proposing their variance to guide interpretation. Nevertheless, we believe there is still considerable room for improvement and exploration in this area. Moreover, we have considered here mixed-memberships, but in certain data-scarce scenarios hard-membership approaches with fewer parameters could be better suited. Future work should consider how to flexibly drive parameters’ inference towards mixed or hard memberships, based on the input data. Our method could be further extended to accommodate distinct community-covariate contributions by integrating two separate $\boldsymbol{H}$ matrices for both in-coming and out-going communities, respectively. This modification will offer clearer insights into how covariates influence the partitions, especially when discrepancies arise between in-coming and out-going communities. Moreover, we treated the number of communities $K$ as given, and for real-world data it was selected via cross-validation. While this procedure is grounded, it can be computationally expensive. An alternative approach would be to treat $K$ as a model parameter and infer it directly from the data. Lastly, it would be interesting to expand this framework to incorporate higher-order interactions, an emerging area that has shown relevance in describing real-world data [44].
@@
Fig. S5: Prediction and community detection performance on synthetic data. We analyze synthetic attributed multilayer networks with $N=500$ nodes, $L=2$ layers (one being assortative and the other disassortative), a categorical attribute with $Z=3$ categories, $K=3$ overlapping communities, and increasing average degrees $\langle{k}\rangle\in\{10,15,20,\dots,50\}$. The results represent averages and confidence intervals over $20$ independent samples. For prediction tasks, we employ a $5$-fold cross-validation procedure. The evaluation metrics include (A) the AUC for edge prediction, with a baseline of $0.5$ corresponding to random choice, and (B) accuracy for covariate prediction. Here, MRF represents a baseline given by the predictions obtained from the maximum frequency in the training set, while Random denotes the uniform random probability over $Z$. (C) Community detection performance is assessed using Cosine Similarity (CS). As inferred point estimates, we consider both the mean of transformed Dirichlet posterior distributions and the $\operatorname{softmax}$ transformation of $\hat{\boldsymbol{\mu}}^{\boldsymbol{\theta}}$. Overall, PIHAM exhibits comparable performance to MTCOV across all tasks despite its broader framework, especially in scenarios involving denser networks.